OpenStudy (anonymous):

What are the solutions of 6x^2 − x − 15 = 0?

OpenStudy (lgbasallote):

just solve for x...factor 6x^2 - x - 15 first

OpenStudy (lgbasallote):

what do you get?

OpenStudy (anonymous):

nothing yet, but thx again for helping

OpenStudy (lgbasallote):

well just factor it then you'll get something like (x+a)(x-b) you just equate the factors to zero x + a = 0 so x = -a x - b = 0 so x = b just like that

OpenStudy (anonymous):

so.... (x+6)(x+1)?

OpenStudy (lgbasallote):

uhmm no...6x 1 is not -15 remember that when you FOIL that you should get your orig problem

OpenStudy (anonymous):

-2.5x6 does, so does that mean that first x is -3/2?

OpenStudy (anonymous):

the second would be either 5/3 or -5/3

OpenStudy (anonymous):

5/3

OpenStudy (anonymous):

i think i understand now....thanks again...

OpenStudy (lgbasallote):

-2.5 x 6? hmmm i would just do it as (3x-5)(2x+3) 3x = 5 therefore x = 5/3 2x = 3 therefore x = -3/2 you were close

OpenStudy (lgbasallote):

maybe your method could work too